OpenStudy (anonymous):

A physics teacher swings a pail of water in a vertical circle 5.0 ft. in a radius at constant speed. (a) What is the speed of the pail of water? (b) What is the maximum time per revolution if the water is not to spill?

OpenStudy (anonymous):

At the top of the circle, gravity must act completely as a centripetal force.\[\underbrace{\cancel mg}_{\text{gravity}}=\underbrace{\frac{\cancel mv^2}{r}}_{\text{centripetal}}\Rightarrow \boxed{v=\sqrt{rg}}\]Since constant velocity is defined by \(v=\frac{\Delta x}{\Delta t}\), we can say the following based on what we know about the geometry of the system. We can use our found value for \(v\) to find \(\Delta t\). \[v=\frac{2\pi r}{\Delta t} \Rightarrow \boxed{\displaystyle \Delta t = 2\pi \sqrt{\frac{r}{g}}}\]

OpenStudy (anonymous):

Again, plug in your given values.